$SPAD/src/input d001aqf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:Expression Float:=(X**2+0.01**2)**-1
f:ASP1(G):=retract e
a:SF:=-1.0
b:SF:=1.0
c:SF:=0.5
epsabs:SF:=0.0
epsrel:SF:=1.0e-4
liw:=200
lw:=4*200
result:=d01aqf(a,b,c,epsabs,epsrel,lw,liw,-1,f)

———-

References^{[1] nothing}

4